OpenStudy (anonymous):

*******************************MEDALS***********************************************(8.6 x 10^8) by (3.2 x 10^3).

OpenStudy (cggurumanjunath):

8.6*3.2 = ?

OpenStudy (cggurumanjunath):

27.52 * 10^(8+3) = 27.52 * 10^11

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html ,go here copy see it work out for answer steps

OpenStudy (anonymous):

nope math papa don't helps

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html , yes does show

OpenStudy (cggurumanjunath):